Calculating Interest Using 360 days as the denominator, calculate interest for the following notes using the formula 1 = P x R x T. If required, round your answers to the nearest cent. Time 30 days 60 Principal Rate Interest $5,400 1,000 3,700 950 1,250 2,100 6.00% 7.50 8.00 6.80 7.25 7.00 120 95 102 90
Answer #1

Formula used as per question is

Interest = Principal x Rate x Terms - days/360

Principal Rate Terms days Interest Calculation
             5,400 6.00% 30            27.00 =5,400*6%*30/360
             1,000 7.50% 60            12.50 =1,000*7.5%*60/360
             3,700 8.00% 120            98.67 =3,700*8%*120/360
                 950 6.80% 95            17.05 =950*6.8%*95/360
             1,250 7.25% 102            25.68 =1,250*7.25%*102/360
             2,100 7.00% 90            36.75 =2,100*7%*90/360